### v1

@@ Abstract
As large language models (LLMs) become increasingly central to Arabic NLP applications, evaluating their understanding of regional dialects and cultural nuances is essential, particularly in linguistically diverse settings like Saudi Arabia. This paper introduces Absher, a comprehensive benchmark specifically designed to assess LLMs performance across major Saudi dialects. Absher comprises over 18,000 multiple-choice questions spanning six distinct categories: Meaning, True/False, Fill-in-the-Blank, Contextual Usage, Cultural Interpretation, and Location Recognition. These questions are derived from a curated dataset of dialectal words, phrases, and proverbs sourced from various regions of Saudi Arabia. We evaluate several state-of-the-art LLMs, including multilingual and Arabic-specific models. We also provide detailed insights into their capabilities and limitations. Our results reveal notable performance gaps, particularly in tasks requiring cultural inference or contextual understanding. Our findings highlight the urgent need for dialect-aware training and culturally aligned evaluation methodologies to improve LLMs performance in real-world Arabic applications.
@@
Large Language Models, Natural Language Processing , Dialects Understanding
@@ 1 Introduction
Recent breakthroughs in the area of language processing have shown that contemporary computational models can perform a variety of tasks, including text production, question answering, and translation between languages [1, 2]. Most of the models, though, have been developed with a prevailing focus on English and Modern Standard Arabic (MSA), limiting their performance in domains with linguistic and cultural diversity, such as Saudi Arabia. These systems tend to find it difficult to interpret expressions related to local identity like idiomatic expressions, cultural practices, and dialectal tones that are underrepresented in the data used to train them [3, 4].
Saudi Arabia, as depicted in Figure 1, have a unique and rich linguistic landscape characterized by a significant variety of regional dialects. These include Central, Western, Southern, Eastern, and Northern, each with distinct features in pronunciation, vocabulary, and grammar [5]. Beyond phonetic and lexical differences, these dialects are intimately linked to region-specific proverbs, traditional expressions, and patterns of social interaction, forming a fundamental part of ordinary conversation. While essential, these features are rarely incorporated into existing language resources or evaluation tools [6, 7]. Therefore, there is growing academic interest in exploring ways to enable the proper interpretation of such dialectal and cultural features in computational models.
Figure 1: A map of Saudi Arabia showing the different regions and their dialectical variations.
Beyond improving technical accuracy, interpreting Saudi dialects also has broader sociolinguistic significance. These dialects are not only linguistic systems but also carriers of regional identity, social norms, and cultural heritage. Ensuring that LLMs can process and respond to them fairly contributes to linguistic inclusivity and mitigates the marginalization of underrepresented communities. Such efforts are vital for the development of equitable AI systems and for informing culturally aware language technologies and public policy.
Current benchmarks for Arabic language models are based primarily on MSA or generalized applications [8, 9], with very little attention paid to dialect and cultural representation. This under-representation affects the fairness as well as the effectiveness of LLMs when utilized by Arabic speakers in real-world situations. In particular, Saudi dialects and cultural elements are less represented in training sets and test sets, leading to inferior performance on culturally contextualized tasks like the interpretation of idiomatic expressions or answering appropriately in contextually sensitive contexts [10, 11].
@@
To address this critical gap, our work introduces Absher, a novel benchmark specifically designed to assess the performance of large language models (LLMs) on the unique and fine-grained linguistic and cultural nuances of Saudi dialects. As illustrated in Figure 2, Absher provides a comprehensive evaluation framework built upon a curated set of dialectal contents—including words, phrases, and proverbs—representing diverse regions across Saudi Arabia. The benchmark comprises a total of 18,564 multiple-choice questions, systematically generated to cover six distinct types: Meaning, True/False, Fill-in-the-Blank, Contextual Usage, Cultural Interpretation, and Location Recognition.
Figure 2: Overview of our benchmark.
In our work, we aim to provide answers to these key questions. First, we investigate the ability of LLMs to interpret Saudi dialects and cultural expressions by creating a new benchmark dataset that captures the nuances of both. Second, we identify the main challenges LLMs face in this context through a detailed error analysis of their performance on this benchmark. Finally, we present a comparative analysis of different LLMs, evaluating their strengths and weaknesses in handling Saudi dialect and culture-specific tasks.
Our work addresses the aforementioned research questions through the following core objectives:
  • Benchmark Development: Construct a culturally grounded benchmark that systematically evaluates LLMs across diverse Saudi dialects. The benchmark includes multiple question formats targeting both linguistic comprehension and cultural awareness, enabling fine-grained assessment.
  • Performance Evaluation: Conduct a rigorous evaluation of several LLMs using the constructed benchmark to assess their capabilities in interpreting Saudi dialects. The analysis identifies their strengths, limitations, and areas that require further research and refinement.
@@ 2.1 Arabic Dialects in LLMs
Recent advances in Arabic NLP have led to increased attention toward evaluating large language models (LLMs) on dialectal and cultural understanding. Several benchmarks have been proposed to assess model performance in these dimensions. For instance, AraDiCE [7] introduced a benchmark targeting dialectal and cultural understanding across Levantine, Egyptian, and Gulf Arabic, revealing critical limitations in generation, translation, and reasoning capabilities. Calibration-Arabic-PLMs [12] focused on model confidence calibration in dialect identification tasks, highlighting inconsistencies between model certainty and actual performance. Health-LLM-Arabic [13] evaluated LLMs like GPT-4 on health-related claims involving Saudi, Egyptian, and Moroccan dialects, and proposed the Cultural Sensitivity Score to quantify dialect adaptation in clinical scenarios.
Complementing these evaluation-focused efforts, a number of datasets and models have been introduced to enhance dialect and culture awareness in Arabic LLMs. Dallah [14] presented a dialect-aware multimodal assistant fine-tuned across six Arabic dialects. AraReasoner [15] provided a benchmarking suite for assessing the reasoning capabilities of LLMs across a wide range of Arabic NLP tasks. Jawaher [16] curated a multidialectal dataset of Arabic proverbs designed to test LLMs’ understanding of figurative and culturally grounded expressions. Similarly, CARE [17] introduced a multilingual benchmark with human preference ratings aimed at evaluating cultural alignment in Arabic and Chinese contexts. Palm [18] contributed a linguistically and culturally diverse dataset spanning 22 Arab countries, with instructions formulated in both MSA and regional dialects to support comprehensive evaluation.
In the multimodal space, several resources have emerged to test vision-language capabilities in culturally rich Arabic settings. Peacock [19] introduced a family of Arabic multimodal LLMs and released Henna, a benchmark designed to assess cultural competence. JEEM [20] focused on visual question answering and image captioning across four Arabic dialects, using culturally specific imagery. Lastly, Swan and ArabicMTEB [21] offered a family of Arabic-centric embedding models and a robust benchmark suite for evaluating cross-lingual, multi-dialectal, and multi-domain embedding quality.
@@ 2.2 Saudi Dialects in LLMs
Despite growing interest in Arabic NLP, few studies have explicitly focused on evaluating large language models (LLMs) within the context of Saudi dialects. One of the first targeted efforts is SaudiCulture [6], a benchmark designed to assess LLMs’ cultural competence across five major regions of Saudi Arabia. Evaluations of models such as GPT-4, LLaMA 3.3, FANAR, Jais, and AceGPT revealed notable performance degradation when addressing region-specific or culturally grounded questions.
Similarly, AL-QASIDA [22] evaluated nine LLMs across eight Arabic dialects, including Saudi Arabic, analyzing generation quality through fidelity, fluency, and diglossia metrics. The study highlighted the impact of pre-training data and suggested methods for dialect-aware model adaptation. In broader benchmarks such as Dallah [14] and Palm [18], Saudi dialects were included alongside other regional variants, though without a focused or fine-grained analysis of Saudi-specific linguistic features.
Beyond direct evaluation, other efforts have explored data creation and translation to support dialectal Arabic processing. For instance, LLM-MT-Data [23] employed LLMs to translate dialectal Arabic—sourced from the Saudi Audio Dataset for Arabic (SADA)—into Modern Standard Arabic, demonstrating the potential of LLMs in generating high-quality parallel data. Similarly, AraBench [24] tackled dialect-specific machine translation using fine-tuning and back-translation techniques to improve Arabic–English translation quality.
While these initiatives mark important progress, most prior work has either concentrated on broader Arabic dialects or MSA, without a dedicated focus on the linguistic and cultural intricacies of Saudi dialects. Dialects such as Central, Western, and Southern are deeply intertwined with region-specific expressions, proverbs, and cultural practices that remain underexplored in current benchmarks. Notably, no prior study has systematically examined Saudi dialects across diverse linguistic units—such as words, phrases, and proverbs—or assessed LLMs’ ability to interpret them in context. This gap underscores the need for a benchmark like Absher, which provides a structured, culturally grounded evaluation of LLMs within the Saudi dialect landscape.
@@ 3 Constructing Absher Benchmark
To assess the ability of computational systems to understand the dialectal and cultural nuances specific to Saudi Arabia, we introduce the Absher Benchmark. This evaluation framework is built upon a curated set of dialectal contents—including words, phrases, and proverbs—representing diverse regions across Saudi Arabia. The development process followed a structured pipeline encompassing data collection, task formulation, prompt engineering, automated question generation, and expert-based review as seen in Figure  3.
@@ 3.1 Data Collection
To evaluate the ability of computational models to interpret Saudi dialects and culturally embedded expressions, we introduce Absher, a benchmark grounded in authentic regional content. It is built from a curated collection of dialectal items—including words, phrases, and proverbs—representing linguistic diversity across Saudi Arabia. The benchmark construction followed a structured pipeline encompassing data collection, content classification, task formulation, prompt engineering, automated question generation, and expert validation, as illustrated in Figure 3.
The core of the dataset was collected from the Moajam website [^1] https://ar.mo3jam.com/, a digital repository that aggregates dialectal vocabulary across Arabic regions. We scraped and organized entries into three categories: words, phrases, and proverbs, each annotated with its meaning, dialect label, contextual usage, and translation. This process ensured comprehensive coverage of both literal and figurative expressions relevant to regional Saudi Arabic.
Words refer to vocabulary items specific to Saudi dialects, often including slang or regional variants not found in Modern Standard Arabic—for example, ”Rabshah”.
Phrases encompass idiomatic or conversational expressions commonly used across regions, such as ”Allah Yta’anni”.
Proverbs are culturally rooted sayings that convey implicit meanings, requiring contextual and cultural understanding—e.g., ”Ally ma ya’arf alsger yshwaeh”.
The raw collection initially consisted of 5,483 items, which were refined through preprocessing and filtering, resulting in 3,094 unique contents used to generate over 18,000 multiple-choice questions across six task types. Table 1 summarizes the dataset statistics.
@@
Table 1: Content statistics before and after preprocessing and the number of generated questions by category.
@@ 3.3 Prompt Design
To rigorously assess the capabilities of LLMs in interpreting Saudi dialects, we employed a structured prompt design strategy tailored to dialectal understanding. Each prompt template included four core components: a task description, a dialect-based question, multiple-choice answer options, and a designated correct answer (see Table 2). Prompts were derived from dialectal contents—words, phrases, and proverbs—each annotated with its dialect label, meaning, and a contextual usage example.
The initial prompt generation followed a simple formulation process, but several issues emerged. Correct answers frequently appeared in fixed positions (e.g., always option A), True/False questions were heavily biased toward “True,” and some questions included incorrect dialectal labels. These inconsistencies introduced unintended biases, undermining the fairness and reliability of the benchmark.
To address these limitations, we refined the prompt design by introducing structural controls. Correct answer positions were randomized to prevent positional bias, and the distribution of True/False items was balanced to ensure fairness. Dialect annotations were reviewed and validated by linguistic experts to confirm their accuracy. These refinements significantly improved the robustness, validity, and representativeness of the final set of multiple-choice questions, establishing Absher as a dependable tool for dialect-sensitive LLMs evaluation. For the complete prompt template, please refer to Appendix A.

### v2

@@ Abstract
As large language models (LLMs) become increasingly central to Arabic NLP applications, their effectiveness in linguistically diverse settings, particularly regions with rich dialectal variation such as Saudi Arabia, remains underexplored. Existing evaluation paradigms tend to prioritize high-resource languages or Modern Standard Arabic (MSA), overlooking regional linguistic and cultural specificities. This leads to performance limitations and cultural biases in real-world deployments. To address this gap, we introduce Absher, the first comprehensive and fine-grained benchmark designed to assess the understanding of LLMs regarding Saudi dialects and their embedded cultural nuances. Absher consists of over 18,000 multiple choice questions derived from a curated dataset of dialectal words, phrases, and proverbs sourced from five major Saudi regions. The benchmark spans six task categories: Meaning, True/False, Fill-in-the-Blank, Contextual Usage, Cultural Interpretation, and Location Recognition, enabling multifaceted evaluation across both linguistic and cultural dimensions.We perform zero-shot evaluations on six state-of-the-art open LLMs: ALLaM, LLaMA, Jais, Mistral, Qwen, and AceGPT. Our results reveal substantial performance variability across dialects and question types. Qwen achieved the highest overall accuracy, excelling in word-level questions (63%), while ALLaM outperformed others in the interpretation of proverbs (48% accuracy). All models struggled with content from underrepresented dialects, particularly Southern and Eastern variants, and with context-free True/False questions, highlighting weaknesses in dialect grounding and binary reasoning. These findings demonstrate the need for dialect-aware training and culturally aligned evaluation. We position Absher as a critical step towards more equitable and effective LLM development for real-world Arabic applications.
@@
Large Language Models, Natural Language Processing , Cultural aware LLMs, Dialects Understanding
@@ 1 Introduction
Recent breakthroughs in the area of language processing have shown that contemporary computational models can perform a variety of tasks, including text production, question answering, and machine translation [Brown2020, OpenAI2023]. These powerful computational models are becoming indispensable tools across various sectors, from education and healthcare to public services and entertainment. However, a critical challenge remains: the vast majority of these models, along with their foundational training data and evaluation benchmarks, are developed with a prevailing focus on high-resource languages like English or, within the Arabic-speaking world, Modern Standard Arabic (MSA). This limits their performance and applicability in domains characterized by significant linguistic and cultural diversity, particularly those with rich regional dialects [Baly2023, alwajih2025palm]. Language is inherently cultural; its full comprehension demands reasoning beyond literal meaning to grasp social implications and regional specificities.
Saudi Arabia exemplifies this challenge with its rich landscape of regional dialects. These dialects are not merely linguistic variants, but embody unique cultural practices, idiomatic expressions, and proverbs critical to authentic communication [al2018suar]. Despite this profound linguistic and cultural diversity, existing Arabic natural language processing (NLP) benchmarks and large language models (LLMs) training data largely overlook these fine-grained Saudi dialectal and cultural nuances [ayash2025saudiculture, Baly2023]. Therefore, models struggle to interpret local identity, leading to significant performance gaps and potentially biases against underrepresented languages.
@@
To address this critical gap and rigorously assess LLMs’ cultural and linguistic competence in this domain, we introduce Absher, a novel, comprehensive, and first-of-its-kind benchmark. Absher offers an evaluation framework built on a curated dataset of authentic Saudi dialectal content, sourced from diverse regions across the Kingdom. It comprises $18,564$ multiple-choice questions, systematically generated to cover six distinct categories: Meaning, True/False, Fill-in-the-Blank, Contextual Usage, Cultural Interpretation, and Location Recognition. This multi-faceted design facilitates a deep, granular assessment of LLMs’ capabilities that moves beyond superficial linguistic understanding toward genuine cultural competence.
Beyond technical accuracy, the ability of LLMs to process and respond to Saudi dialects carries broader sociolinguistic significance. These dialects are not only linguistic systems but also carriers of regional identity, social norms, and cultural heritage. Ensuring their equitable representation in AI systems is crucial for linguistic inclusion and mitigating the marginalization of underrepresented communities. Such efforts are vital for the development of equitable AI systems and for informing culturally aware language technologies and public policy.
Figure 1: Overview of Absher, covering all regions of Saudi Arabia (five specific regions and a general category) and presents representative examples from its three types of content: words, phrases, and proverbs. The benchmark also spans all six question types: Meaning, True/False, Fill-in-the-Blank, Contextual Usage, Cultural Interpretation, and Location Recognition.
In this work, our objective is to provide answers to key research questions concerning LLMs’ capabilities in interpreting Saudi dialects, the specific challenges they face, and a comparative analysis of their performance. Our goal is twofold: (1) to develop a novel, fine-grained benchmark that systematically evaluates LLMs across Saudi dialects and linguistic units; and (2) to conduct a comprehensive evaluation of state-of-the-art LLMs, analyzing their performance, weaknesses, and areas for improvement. Our key contributions are as follows:
  • Benchmark Development: We introduce Absher, the first large-scale benchmark tailored to Saudi dialects, featuring over 18,000 multiple-choice questions. These questions span six task categories and are grounded in authentic dialectal words, phrases, and proverbs from five major Saudi regions, addressing a critical gap in Arabic NLP evaluation. Figure  1 illustrates the composition of Absher across different regions and categories.
  • Methodological Framework: We propose a scalable pipeline that combines prompt engineering with automated question generation using LLMs, followed by a rigorous multi-stage human validation process. This ensures high linguistic and cultural fidelity, making Absher a reliable tool for dialectal evaluation.
  • Model Evaluation and Analysis: We conduct a comprehensive evaluation of multilingual and Arabic-specific LLMs on the Absher benchmark. Our analysis reveals significant performance variability across dialects, task categories, and linguistic units. We also investigate the impact of model size, language orientation, and architecture, offering insights into the limitations and strengths of current LLMs in dialect-sensitive settings.
This paper is structured as follows: Section 2 reviews the existing literature on Arabic dialects in LLMs. Section 3 details the methodology for constructing the Absher benchmark, including data collection, pre-processing, prompt design, question generation, and human evaluation. Section 4 presents the experimental setup and results of our LLM evaluations. Section 5 provides a detailed analysis of our findings, including comparative performance across models, content types, and question types. Finally, Section 6 discusses the broader implications of our results and outlines promising directions for future research.
@@ 2.1 Arabic Dialects in LLMs
Recent efforts in Arabic NLP have increasingly emphasized dialectal and cultural aspects in evaluating LLMs. Several benchmarks have emerged to measure model performance in these dimensions. For example, AraDiCE [Baly2023] assessed LLMs on dialectal and cultural understanding across Levantine, Egyptian, and Gulf Arabic, highlighting limitations in generation and reasoning. Calibration-Arabic-PLMs [Al-Laith2025] examined model confidence in dialect identification, revealing misalignment between confidence and accuracy. Health-LLM-Arabic [alharbi-etal-2025-evaluating] introduced a Cultural Sensitivity Score to evaluate model responses to health-related claims in Saudi, Egyptian, and Moroccan dialects.
In addition, to these evaluation efforts, several datasets and models have been designed to enhance dialect awareness. Dallah [alwajih2024dallah] developed a multimodal assistant fine-tuned on six dialects, while AraReasoner [hasanaath2025arareasoner] proposed a benchmark suite for evaluating Arabic reasoning capabilities. Jawaher [magdy2025jawaher] focused on figurative language through a collection of Arabic proverbs, and CARE [guo2025care] evaluated cultural alignment via human preferences in Arabic and Chinese contexts. Palm [alwajih2025palm] contributed a linguistically diverse dataset covering 22 Arab countries, combining MSA and dialectal instructions.
In the multimodal space, Peacock [alwajih2024peacock] released the Henna benchmark to test cultural competence, and JEEM [kadaoui2025jeem] introduced dialect-aware visual tasks involving culturally grounded images. ArabicMTEB and Swan [bhatia2024swan] offered resources for evaluating cross-lingual, dialectal, and domain-specific Arabic embeddings.
@@ 2.2 Saudi Dialects in LLMs
Despite growing interest in Arabic dialects, Saudi-specific evaluations remain limited. SaudiCulture [ayash2025saudiculture] evaluated five LLMs across five Saudi regions, revealing performance drops on culturally grounded content. AL-QASIDA [robinson2024qasida] examined nine LLMs across eight dialects—including Saudi Arabic—using fidelity and diglossia metrics, and discussed pretraining implications. Other benchmarks, such as Dallah [alwajih2024dallah] and Palm [alwajih2025palm], included Saudi dialects but lacked focused analysis. LLM-MT-Data [Elmogtaba2024] used LLMs to translate Saudi speech data into MSA, and AraBench [sajjad2020arabench] explored dialect-to-English translation with fine-tuning and back-translation.
A notable data-centric effort is the KIND dataset [yamani2024kind], which collected fine-grained dialectal Arabic through a community-driven competition. KIND includes parallel corpora and Q&A spanning 29 Saudi dialects, with city-level granularity. However, it is primarily designed for data collection and does not offer a structured evaluation of LLMs’ contextual understanding, a gap Absher seeks to fill.
A key point of contrast is with SaudiCulture [ayash2025saudiculture], which introduced a pioneering cultural benchmark. However, it is limited to 441 questions focused mainly on general knowledge and cultural themes predominantly in MSA. In contrast, Absher provides a broader and more targeted evaluation with a significantly larger volume of questions. It is uniquely designed for a fine-grained, dialect-sensitive assessment across Saudi dialects, focusing on the nuanced understanding of words, phrases, and proverbs. Unlike SaudiCulture, which emphasizes high-level cultural reasoning, Absher enables a deep, multi-category evaluation drawn from region-specific vernacular data, making it a more comprehensive and regionally grounded resource for advancing Arabic NLP.
While these studies offer valuable insights, none provide a comprehensive benchmark that systematically evaluates LLMs on Saudi dialects across linguistic units such as words, phrases, and proverbs. Existing work often overlooks the sociocultural specificity of regional expressions. In contrast, Absher introduces a fine-grained, culturally contextualized benchmark that spans multiple Saudi dialects and domains, enabling detailed evaluation of LLMs’ dialectal and cultural understanding. It further differentiates itself through real-world question design, multi-unit coverage, and region-aware analysis of model behavior.
@@ 3 Constructing Absher Benchmark
To assess the ability of computational systems to understand the dialectal and cultural nuances specific to Saudi Arabia, we introduce the Absher Benchmark. This evaluation framework is built upon a curated set of dialectal contents, including words, phrases, and proverbs, representing diverse regions across Saudi Arabia. The development process followed a structured pipeline encompassing data collection, pre-processing, prompt engineering, automated question generation, and expert-based review as seen in Figure 2.
@@ 3.1 Data Collection
To evaluate the ability of computational models to interpret Saudi dialects and culturally embedded expressions, we introduce Absher, a benchmark grounded in authentic regional content. It is built from a curated collection of dialectal items, including words, phrases, and proverbs, representing linguistic diversity across Saudi Arabia.
The core of the dataset was collected from the Moajam website [^1] https://ar.mo3jam.com/, a digital repository that aggregates dialectal vocabulary across Arabic regions. We scraped and organized entries into three categories: words, phrases, and proverbs, each annotated with its meaning, dialect label, contextual usage, and translation. This process ensured a comprehensive coverage of both literal and figurative expressions relevant to regional Saudi Arabic.
  • Words refer to vocabulary items specific to Saudi dialects, often including slang or regional variants not found in Modern Standard Arabic.
  • Phrases encompass idiomatic or conversational expressions commonly used across regions.
  • Proverbs are culturally rooted sayings that convey implicit meanings, requiring contextual and cultural understanding.
The raw collection initially consisted of 5,483 items, which were refined through pre-processing and filtering, resulting in 3,094 unique contents used to generate over 18,000 multiple-choice questions across six task types. Table 1 summarizes the dataset statistics.
Table 1: Content statistics before and after pre-processing and the number of generated questions by category.
@@ 3.3 Prompt Design
To rigorously assess the capabilities of LLMs in interpreting Saudi dialects, we employed a structured prompt design strategy tailored to dialectal understanding. Each prompt template included four core components: a task description, a dialect-based question, multiple-choice answer options, and a designated correct answer. Table 2 provides an illustration. The prompts were derived from dialectal contents, such as words, phrases, and proverbs. Each annotated with its dialect label, meaning, and a contextual usage example.
The initial prompt generation followed a simple formulation process, but several issues emerged. Correct answers frequently appeared in fixed positions (e.g., always option A), True/False questions were heavily biased toward “True”, and some questions included incorrect dialectal labels. These inconsistencies introduced unintended biases, undermining the fairness and reliability of the benchmark.
To address these limitations, we refined the prompt design by introducing structural controls. The correct answer positions were randomized to prevent positional bias, and the distribution of True/False items was balanced to ensure fairness. Dialect annotations were reviewed and validated by linguistic experts to confirm their accuracy. Crucially, we implemented a rigorous two-step filtering strategy to handle abnormal model outputs from GPT-4o:
  • Automated Filtering (Formatting and Structure): All generated outputs were subjected to a validation script to check for strict structural integrity. This script automatically flagged and discarded any questions that contained malformed JSON output, failed to include exactly four distinct multiple-choice options, or did not designate a clear correct answer. This ensured only well-structured, testable items proceeded to the next phase.
  • Manual Curation (Hallucination and Semantic Errors): A team of two native Saudi Arabic speakers and linguistic experts manually reviewed a sample of automatically validated questions. Their primary task was to filter out semantic errors and hallucinations, specifically removing questions where the generated options were culturally inaccurate, the question’s premise contradicted the source material (meaning or context), or the designated correct answer was objectively wrong.
These refinements significantly improved the robustness, validity, and representativeness of the final set of multiple-choice questions, establishing Absher as a dependable tool for dialect-sensitive LLMs evaluation. For the complete prompt template, please refer to Appendix A.
Table 2: Template for question formulation